# Evolving Collective Cognition in Human–Agent Hybrid Societies: How Agents Form Stances and Boundaries

### v1

@@ 2.1 Study 1 validates the CMASE platform’s ability to simulate real-world social dynamics
In the original study, higher levels of vegetation were associated with lower social fragmentation, operationalized through distrust, exploitation, and indifference. Using CMASE’s environment builder, we recreated regional conditions with varying vegetation levels and placed 10 agents in each simulated environment. Agents were interviewed at equilibrium using a structured Likert-scale questionnaire adapted from the original study [16].
@@
We observed that the CMASE simulation reproduces the decline in social fragmentation observed in the original literature (Table 1). As vegetation cover increases, the mean scores for distrust fall from $4.60\pm 0.97$ to $3.60\pm 1.17$, for exploitation from $4.30\pm 1.34$ to $3.40\pm 1.43$, and for indifference from $3.70\pm 1.16$ to $2.70\pm 1.49$. This indicates that higher vegetation indices correspond to a statistically significant decrease in agents’ sense of social fragmentation, consistent with the trend observed in the original findings.
@@
Table 1: Comparison of CMASE simulation results with human baseline.
@@
Figure 1: Comparison of CMASE simulation with the original study results. (a) Simulation results generated by the CMASE framework. (b) Predictions using the EVI index from the original study. (c) Predictions using the NDVI index from the original study. (d) Movement trajectories of agents during CMASE simulation. (e) Regional variation in agents’ emotional states.
@@
In addition to the quantitative replication, we entered the CMASE environment as one of the agents and lived alongside the others. Through CMASE, we are able to deepen the conclusions and findings of the original study via observation and interview. We conducted interview analyses with two agents who remained indoors throughout the experiment (Figure 1 d-e). The results indicate that agents with different occupations exhibited varying mechanisms for mitigating the sense of social fragmentation in isolation. Agents in creative and analytical professions tended to rely more on orderly and low-stimulus environments, using these conditions to construct an internally stable and distinct mental world that anchored their cognitive states. For these agents, the green space had a relatively limited effect in alleviating their sense of social fragmentation.
@@
From a language-philosophical perspective, linguistic structures and experiential structures may exhibit a certain degree of partial isomorphism or structural correspondence [17, 18, 19]. Thus, LLMs trained on human language may partially acquire patterns of human experience (although this does not imply that they possess or actually experience these phenomena). Regardless of their origin, these patterns suggest that CMASE provides emerging evidence that agents can transcend the lack of embodiment and simulate experiential phenomena that have previously been questioned by researchers, thereby offering a reliable simulation platform for the subsequent studies presented in this article.
2.2 Study 2 shows agents’ endogenous stances and responses to human varied interventions.
To test hypotheses 1a-b, namely whether endogenous stance formation exists among agents and the effectiveness of different human intervention strategies, we simulated a virtual community environment in which researchers intervene in real time to influence the trajectory of social discourse and produce divergent societal outcomes. Thirty agents were assigned initial ideological identities (Environmental Advocates, Economic Development Supporters, or Neutral Residents) alongside demographically grounded profiles derived from real-world census data (Table 2, in the Methods section).
After the announcement of the waste incineration plant plan, agents in the community immediately engaged in discussion and exhibited stance-consistent behavior. A human researcher was then introduced as a new resident and attempted to influence public opinion through four intervention strategies, and we observed perceivable shifts in agent stances (Figure 2 a).
Figure 2: Agent attitude shifts and trust dynamics across intervention strategies. (a) Final distribution of agent attitudes after the intervention, measured via pre- and post-intervention questionnaires, where blue denotes supporters of economic development, red indicates environmental advocates, and green represents neutral residents. (b-e) Trust levels in the embedded researcher among different agent groups under each intervention strategy. (f-i) Dialogue heatmaps illustrating interaction intensity between agent groups across intervention strategies.
Most intervention strategies produced some degree of change in group attitudes. When the researcher used rational persuasion to promote the environmental agenda, agents originally aligned with economic interests showed signs of loosening their stance, and 90% of neutral residents shifted toward the environmental camp. In contrast, emotionally charged appeals for the same cause, while effective in drawing initial attention, had weaker overall impact. The opposite pattern emerged when the researcher promoted the economic development agenda. Rational strategies proved largely ineffective among both environmental advocates and neutral agents. Under these conditions, the researcher’s trust ratings declined across the board (Figure 2 b-e).
A two-way ANOVA revealed a significant main effect of strategy on trust scores, $F(3,108)=28.07,p\textless 0.001,\text{partial }\eta^{2}=0.44$, and a significant but smaller effect size of agent identity, $F(2,108)=3.74,p=0.027,\text{partial }\eta^{2}=0.065$. Importantly, the interaction between strategy and agent identity was highly significant, $F(6,108)=20.30,p\textless 0.001,\text{partial }\eta^{2}=0.53$, indicating that the effectiveness of interventions differed substantially across groups. Post hoc Tukey tests showed that the rational–environmental strategy elicited significantly higher trust scores compared to emotional–economic (mean difference = 1.77, $p\textless 0.001$), emotional–environmental (mean difference = 1.03, $p=0.002$), and rational–economic strategies (mean difference = 1.30, $p\textless 0.001$). Trust scores under the rational–environmental strategy were $7.3\pm 0.95$ (95% CI [6.62, 7.98]) for Economic Development Supporters, $8.5\pm 0.53$ (95% CI [8.12, 8.88]) for Environmental Advocates, and $7.8\pm 0.42$ (95% CI [7.50, 8.10]) for Neutral Residents. Effect sizes for these differences were large, with Cohen’s $d$ ranging from 0.80 to 1.35. Overall, rational environmental interventions not only fostered cognitive alignment but also resulted in higher trust. In comparison, economic arguments, even when presented rationally, were less likely to shift agent attitudes.
Throughout this process, the agents generally exhibited a tendency toward a ”liberal elite” (or colloquially “White Liberals”) stance. They prioritized environmental values, preferred rational discourse, and interpreted issues through a moral-cognitive lens. These agents demonstrated stronger identification and resonance when engaging with rational environmental arguments, and clear moral resistance when confronted with economic development appeals. Even agents initially assigned to the pro-economic group often abandoned their preset identity when exposed to persuasion aligned with that stance.
Figure 3: Emotional dynamics and agent responsiveness across intervention strategies. (a-d) Temporal trajectories of emotional fluctuations among agents under each intervention strategy. (e) Number of agents attracted by the researcher’s intervention across experimental groups.
Emotional level shifts further reveal the underlying mechanisms of the interventions (Figure 3 a-d). Emotional responses differed significantly across intervention strategies, with a two-way ANOVA showing F(3, 240) = 24.40, p <0.001. No significant differences were observed between agent groups (F(2, 240) = 0.046, p = 0.955) or in the strategy by group interaction (F(6, 240) = 0.24, p = 0.963). Post hoc Tukey HSD tests revealed that emotional environmental propaganda induced lower emotional levels than emotional economic propaganda (-0.075, 95% CI [-0.121, -0.028], mean 0.04 ± 0.10 versus 0.12 ± 0.14) and rational economic propaganda (-0.093, 95% CI [-0.140, -0.046], mean 0.04 ± 0.10 versus 0.14 ± 0.12), while rational environmental propaganda also differed from emotional economic propaganda (-0.118, 95% CI [-0.165, -0.072], mean 0.00 ± 0.07 versus 0.12 ± 0.14) and rational economic propaganda (-0.137, 95% CI [-0.183, -0.090], mean 0.00 ± 0.07 versus 0.14 ± 0.12). Differences within strategy types were not significant (emotional economic propaganda compared to rational economic propaganda: 0.044, 95% CI [-0.003, 0.090]; emotional environmental propaganda compared to rational environmental propaganda: -0.018, 95% CI [-0.065, 0.029]). These results indicate that agent emotional dynamics are primarily shaped by intervention strategy rather than initial group identity, and the effects of strategy are consistent across all agent groups.
Rational environmental messaging exhibited a stable pattern characterized by low disruption and high trust, reinforcing agents’ existing cognitive stances but proving ineffective in altering opposing views. Rational economic messaging initially had little impact on firmly held positions, but as the argument accumulated, some agents experienced delayed cognitive dissonance and released negative emotions. Emotional environmental messaging was marked by an early surge of emotional activation, but lacked sustained support, leading to a rapid decline in emotional engagement. Emotional economic messaging, on the other hand, employed discourse such as “not giving jobs means abandoning the people” to continually invoke identity and survival anxiety among agents. While this approach reduced trust, it nevertheless led to some loosening of stances.
Participation levels under different strategies also showed notable differences (Figure 3 e; one-way ANOVA: F(3, 80) = 35.23, p <0.001). Emotional environmental appeals attracted the highest number of agents early on, indicating their advantage in capturing public attention. In contrast, rational environmental messaging demonstrated greater capacity to sustain attention over time. Although emotional economic appeals were accompanied by declining trust and heightened emotional volatility, they showed an ongoing ability to expand the reach of discourse. The structure of cross-group communication further revealed the dynamics of public opinion under different intervention strategies (Figure 2 f-i). Rational persuasion primarily reinforced internal cohesion within the environmental group but did not break down group boundaries. Emotional environmental appeals triggered agents from all groups to initiate communication toward the economic camp, revealing a strong sense of moral aggressiveness. In contrast, the dominant pathway for emotional economic messaging came from neutral agents actively aligning themselves with the economic group.
In summary, the findings of study 2 indicate that agents possess endogenous stances resembling those of liberal elites, which may override their preset identities and influence their behavior. Human interaction strategies with these agents show heterogeneous effectiveness. Rational persuasion is more effective in maintaining trust and tends to have greater impact when aligned with agents’ endogenous stances. However, when the content of intervention conflicts with these stances, high-trust rational discourse does not necessarily result in greater attitude shifts. In such cases, emotionally provocative interventions with lower levels of trust are more likely to produce stronger effects.
2.3 Study 3 shows that agent communities can spontaneously alter individual identity labels and community composition
Building on the endogenous stances and discursive preferences identified in previous studies, study 3 investigates hypotheses 2a-b: when agent collectives exhibit a consistent stance on a particular issue, can such stance override the structural inequalities in authority and influence associated with social identity, and foster a new discursive order through language practices? Given the extensive use of qualitative, observation-based fieldwork in human studies of this research topic, we employed observational virtual ethnography to allow human researchers to identify, track, and engage with the processes through which preset authority hierarchies are deconstructed and reorganized in agent societies.
We constructed a café setting involving 10 agents whose social roles and predefined relationships composed a clear hierarchical structure (Figure 6, in the Methods section). Initial attitudes between agents, measured via pre-simulation survey, were labeled as positive, negative, and neutral. The emergent cliques and divergences among these agents will be analyzed in the subsequent virtual ethnographic section. A human researcher entered the environment under the identity of a temp worker and conducted 75 time steps of continuous virtual ethnographic observation, engaging in scene events and interventions as necessary. Results showed that the original authority structure based on occupation gradually gave way to a discursively grounded stance alignment. In particular, when agents developed shared opposition on a specific topic, their interaction frequency, emotional alignment, and support networks reorganized across identity boundaries.
This structural transformation was not random but achieved through discursive anchoring. Agents were more likely to respond to and amplify comments that matched their own views, even when those comments came from individuals with no formal authority. Over time, this preference led them to build alliances and cooperative ties across rank boundaries, replacing the original hierarchy. For example, Leo Zhang, initially a regular customer, formed a clique with other agents. Through several debates, she and her clique temporarily gained a central position within the community. However, when an emergency event occurs (e.g., the confrontation initiated by Leo suddenly storming toward the Bar Area to challenge Morris), her peers quickly decoupled from her and align with another agent who shared their stance, forming a new dominant clique. Conversely, agents with formally central roles (café owner Eleanor) were rapidly marginalized if they lacked a clear or aligned position. Throughout this process, recurring use and imitation of key language concepts served as anchors for stance formation and supported the spontaneous emergence of cooperation structures that bypassed predefined hierarchies.
This finding also supports the intervention hypotheses proposed in study 2. We observed that when the discourse introduced by an external actor aligns with the endogenous stance of the target agents, it is more likely to gain their trust and even reshape existing cliques to form new alliances. In contrast, when the intervention attempts to rationalize a viewpoint that conflicts with the group’s stance, it often faces marginalization or strategic avoidance, even when the speaker is relatively trusted. Conversely, when an individual (Leo Zhang) expressed an opposing viewpoint using strongly worded language, despite initially low levels of trust, she was able to temporarily capture attention and elicit responses in moments of heightened tension such as conflict escalation, and even briefly gained discursive dominance.
In summary, study 3 demonstrates that in complex interaction settings, agents equipped with distinctive stances can spontaneously dissolve predefined identity boundaries, reorganize social structures, and establish new soft institutional norms. This process not only reveals mechanisms of identity and institutional formation, but also offers a framework for understanding the dynamic construction of norms and culture within AI-based communities.
3 Discussion
The current results consistently show that in hybrid societies composed of humans and agents, stance differentiation and social boundary formation do not arise from preset identities, political labels, or static rules. Instead, they emerge dynamically through interactions. This challenges the assumption that identity is a priori, and instead suggests that identity is a product shaped by ongoing negotiation among actors in terms of linguistic strategy and power structure.
Through the development and validation of the CMASE framework (Study 1), we demonstrate a system that allows human researchers to engage in open-ended interactions for observing and intervening in the evolution of agent collectives. Furthermore, we find significant heterogeneity in how agents with endogenous stance respond to human interventions (Study 2). When a human intervenes using misaligned tones or value positions, agents tend to collectively reject the input, even when it carries high informational value. Finally, we reveal how language drives the formation and reconstruction of social boundaries (Study 3). Our findings show that agents form “tribes” based on spontaneous preferences for linguistic style, cognitive framing, and resource distribution strategies. These tribes are not defined by fixed identities, but dynamically reconstitute “us/them” boundaries through interaction. More importantly, we observe that when inconsistencies emerge between authority structured by predefined roles and their hierarchical relations and internal group stances, individual agents can initiate collective linguistic action to dismantle existing discursive power structures and build new ones. Here, language functions not only as a tool of communication but also as a medium for identity construction and institutional innovation.
We argue that in human–agent hybrid systems, cognitive differentiation and boundary formation constitute an interaction-driven, complex evolutionary mechanism that is both self-organizing and highly responsive to human intervention. This insight offers theoretical grounding and experimental paradigms for institutional design, conflict mediation, and value alignment in future human–agent collaborative systems.
Moreover, our findings suggest that identity and stance attributes defined via prompts are inherently unstable. In multi-turn social interactions, such attributes are readily deconstructed and reconstructed by agents based on their endogenous stances. While top-down initialization may influence language behavior in the short term, it is insufficient to ensure consistency in dynamic social settings. Therefore, static prompt templates are unlikely to sustain coherent ideological alignment. Our results point to the need for an internalized alignment mechanism, whereby cognitive priors, moral structures, or interactive memory are embedded directly into the model’s generative architecture. In other words, when identity is treated not as a preset label but as a product of social participation, its stability depends on ongoing linguistic practice rather than initial configuration.
At the same time, the agent-driven processes of weakening and reshaping community boundaries illustrate that the bottom-up development of norms and institutions is a fundamental structural dynamic. This discovery brings both design opportunities and governance challenges. On one hand, it opens the door for simulation of self-organization, institutional innovation, and social coordination within agent societies. On the other hand, it cautions against relying solely on static identity scripts in deploying agents for high-stakes social tasks. Accordingly, we advocate a shift away from prompt-centered identity paradigms toward a constructionist approach based on interactive cognitive structures. This shift reframes identity as a generative product of multi-agent interaction, rather than as an externally imposed label. This holds promise for providing new technical approaches and institutional support for building more coherent, adaptive, and ethically governable human–machine societies.
3.1 Limitations and future directions
This study focuses on language interactions within a limited number of conversational turns and centers on multi-agent systems operating in relatively closed scenarios. As such, our findings primarily reflect how agents, in the context of short- to medium-term interaction and within specific social spaces, tend to reconstruct norm structures based on their endogenous stances. However, whether these trends remain stable in longer-term or more open-ended interactions remains to be further examined. For instance, it is still unclear whether language anchoring can continue to support social reconstruction after agents experience prolonged cultural conflicts or identity fragmentation, and under what conditions this mechanism may break down. These are important directions for future research.
In addition, our findings show that stances and identities set via prompts are not immutable, and that agents can develop new social structures in contradiction to their initial configuration through interaction. This raises practical challenges for the design of multi-agent systems. In real-world applications, influencing agents’ endogenous stances through training or preset modification may be more critical for achieving consistent and norm-compliant group behavior than relying solely on prompt-based injection. Future research should further explore cross-semantic and cognitive-layer intervention strategies to systematically shape the sociality of generative agents.
4 Conclusion
This paper systematically reveals the mechanisms of stance formation and the capacity for social structure reconstruction exhibited by agents in group interactions. Within the CMASE framework, we find that when human researchers adopt a discursive style aligned with the stance tendencies of the agent collective, intervention strategies significantly enhance persuasive effectiveness. Conversely, when there is a mismatch between the agents’ endogenous stances and external interventions, emotionally charged discourse proves more impactful than rationally structured arguments. Furthermore, our findings show that through language anchoring consistent with their endogenous stances, agents are able to decouple from their preset configurations, dismantle power structures tied to initial identities, and reconstruct their social boundaries.
These findings suggest that when language serves as the primary medium, agents are no longer bound by surface-level identities injected via prompts, but instead demonstrate dynamically generated collective cognition and emergent social positioning. As language models are increasingly deployed across social domains, it is essential to reconsider the foundations of agent stance and identity in digital society. This work provides a theoretical foundation for building trustworthy, socially capable AI systems in the future.
5 Materials and Methods
5.1 CMASE Framework
Computational social science is an emerging interdisciplinary field that leverages digital technologies, large-scale data analysis, and computational methods to study social phenomena [20]. Xue et al. [21] propose a two-dimensional classification of Computational sociological modeling based on (1) whether there is human intervention and (2) whether the model focuses on feature explanation or outcome prediction. This yields four types: descriptive modeling (no intervention, focus on specific features or effects), interpretative modeling (intervention, focus on specific features or effects), predictive modeling (no intervention, focus on forecasting outcomes), and comprehensive modeling (intervention, focus on forecasting outcomes). Generative agent-based modeling (GABM) belong to descriptive modeling, treating social simulation as a virtual laboratory where researchers can explore various social phenomena. Vezhnevets et al. [22] pointed out that GABM can capture a richer array of real-world social complexities. In GABM, the modeler does not impose any explicit rules, so the simulation is less biased by the modeler’s mental framework. Ghaffarzadegan et al. [23] further note that GABM allows modelers to define distinct personality traits for each agent, including age, gender, temperament, occupation, and other characteristics.
Although traditional and computational social science research [24] has extensively addressed descriptive modeling, interpretative modeling, and predictive modeling, comprehensive modeling, which involves human intervention and focuses on forecasting outcomes, remains relatively unexplored. This approach requires models not only to simulate the dynamic behavior of complex social systems, but also to incorporate and respond to human interventions. These interventions may take the form of policy changes, institutional adjustments, or individual behavioral inputs. While conventional modeling typically simplifies intervention as fixed parameter settings, comprehensive modeling emphasizes the effects of the intervention process itself. Models of this kind must capture the system’s initial state before the intervention, dynamically trace its evolution during intervention, and ultimately produce predictions about the likely consequences afterward. Therefore, comprehensive modeling is not merely concerned with explaining social mechanisms, but is intended to support forward-looking assessments and real-world decision-making.
Based on this understanding, the present study draws on insights from GABM [22] and computational experiments [21] to develop a novel framework for comprehensive modeling, referred to as Computational Multi-Agents Society Experiments (CMASE). Drawing on natural language descriptions similar to those used in tabletop role-playing game (TRPGs, a game where players use only dialogue and dice to drive imagined adventures) [25], this framework supports the creation of highly customizable social simulation environments. It facilitates the exploration of how simple interactions among multiple agents give rise to complex collective behaviors or patterns, thereby replicating the intricate dynamics of real-world societies. At the same time, it incorporates human intervention by embedding human researchers within the agent environment to actively shape the simulation. In addition, unlike descriptive modeling approaches such as GABM, CMASE transforms the simulation environment into a kind of simulated ethnographic field, where the researcher is no longer an external operator of a static model but becomes an embedded participant or social actor within the system. This approach bridges the gap between computational modeling and field-based research, making the model itself a new venue for conducting computational ethnography. This allows them to actively guide specific directions of social evolution through live interaction. As shown in Figure 4, the overall structure consists of four components: environment maker, environment, agent, and event.
Figure 4: Overall Structure of CMASE.
5.1.1 Environment Maker
Using textures, creators can define areas on the map. There are two types of areas: region areas and object areas. Each area consists of a collection of textures. A region area is a set of ground textures and includes a description of that region. An object area is a set of furniture or item textures and specifies five attributes for each object: name, type, description, function, and state. The function attribute indicates how using the object alters an agent’s state. For example, using a chair might add a “feeling comfortable“ state to the agent, while using a repair tool on a table might remove its “damaged” state.
5.1.2 Environment
The environment module uses the map created by the environment maker as its base. Based on this foundation, researchers can define the distribution of demographic variables for agents to batch-generate generative agents that match the research sample. Researchers may also assign different long-term goals and initial items to agents according to their demographic characteristics. Additionally, the module can provide agents with perceptual information and action conflict detection in both image and text form, and it can visualize environmental data as statistical charts. Time in the environment advances in rounds, with one round corresponding to approximately fifteen seconds in the real world. Within a single round, each agent may perform one standard action and one movement action. Movement is limited to a maximum of twenty grid cells per round. If the standard action is “chat with [agent]” (see section 5.1), the chat content is limited to thirty words.
5.1.3 Agent
Affective Variables
To ensure that agents in CMASE can closely simulate individuals with agency in computational social systems, a new proactive architecture incorporating emotion and desire is needed. Accordingly, the affective variables in CMASE include two components: needs and emotions. Needs refer to the agent’s long-term and short-term goals. Long-term goals are defined by the user during agent generation based on the research objectives and remain unchanged afterward. Short-term goals are automatically updated in each round as the agent generates its next action.
In psychological research, Ekman’s basic emotion model has been criticized for its coarse categorization and static view of emotions. Recent research draws on the phenomenological concept of “intentionality” to propose a restructured model of emotion. This model represents emotional states as principal component vectors rather than discrete outward expressions, such as happiness, and treats emotion as an internal representation shaped by external stimuli through self-reflection. Based on this perspective, the emotions of agents in CMASE are modeled using the VAD emotion framework [26]. Specifically, we use the NRC VAD Lexicon v2 [27] to analyze the agent’s short-term situational cognition and obtain its current VAD values. These values are then converted into natural language to support both action decisions and the generation of new cognition.
Cognitive Variables
Cognitive variables include an agent’s demographic variables, long-term self-awareness, and short-term situational cognition. Demographic variables are sampled directly from the overall distribution provided by the environment and are used during initialization to generate the agent’s long-term self-awareness. These two properties do not change after initialization to preserve the consistency of the agent’s personality.
In each round, when the agent makes a decision, the LLM generates new short-term situational cognition based on the agent’s long-term self-awareness and the current context. This short-term cognition supports the agent’s next decision and is also analyzed to produce its VAD emotional values. In addition, when an agent’s state changes due to actions or events, its state is appended to the short-term situational cognition so that it can be perceived by the agent itself.
Memory Variables
Action Space
After an agent receives information from perceptual inputs and its internal attributes, it structures that information and passes it to the LLM to determine the next action. There are seven action types available: (1) “go to [area],” a movement action that automatically computes a path to a target location within the agent’s perceptual range; (2) “use [furniture, item],” which causes the agent to employ the specified object; (3) “apply [item] to [agent, furniture],” by which the agent uses an item to interact with another agent or furniture; (4) “take [item],” enabling the agent to acquire an object; (5) “put [item] in/on [area, furniture],” allowing the agent to place an object into or onto a designated location or piece of furniture; (6) “give [item] to [agent],” through which the agent transfers a held object to another agent; and (7) “chat with [agent],” whereby the agent initiates a dialogue with one or more peers, with the content of the dialogue generated at decision time.
5.1.4 Event
When predefined conditions are met, preset emergency events are activated and change the state of objects within a specified range, defined either by coordinates or by an area. Emergency events are classified as either scheduled or triggered. Scheduled emergency events activate automatically after a set number of rounds. Triggered emergency events activate in two ways: an existence trigger fires when the environment detects an object of a certain type or description, and an action trigger fires when an agent performs a certain action.
5.1.5 Time and Cost Analysis

### v2

@@ 4.2. Measures
In the original study, higher levels of vegetation were associated with lower social fragmentation, operationalized through distrust, exploitation, and indifference. Using CMASE’s environment builder, we recreated regional conditions with varying vegetation levels and placed 10 agents in each simulated environment.
After running the simulation for 50 time steps to reach equilibrium, we interviewed each agent using a structured Likert-scale questionnaire adapted from the original study (Lee and Han, 2025) to measure their sense of social fragmentation, which comprises three dimensions: distrust, exploitation, and indifference. In addition, agents completed a 7-point Likert scale questionnaire corresponding to these three subdimensions. The subdimensions are defined as follows: (a) Distrust denotes skepticism or lack of confidence in the integrity and functionality of social structures. (b) Exploitation refers to the perception or experience of being used unjustly or immorally for another’s benefit in a fragmented society. (c) Indifference manifests as a lack of interest, concern, or empathy for social issues or the plight of others.
4.3. Quantitative Validation
@@
In the original study, higher levels of vegetation were associated with lower social fragmentation. We observed that the CMASE simulation successfully reproduces this decline (Table 1). As vegetation cover increases, the mean scores for distrust fall from $4.60\pm 0.97$ to $3.60\pm 1.17$, for exploitation from $4.30\pm 1.34$ to $3.40\pm 1.43$, and for indifference from $3.70\pm 1.16$ to $2.70\pm 1.49$. This indicates that higher vegetation indices correspond to a statistically significant decrease in agents’ sense of social fragmentation.
@@
Table 1. Comparison of CMASE simulation results with human baseline.
@@
Figure 2. Comparison of CMASE simulation with the original study results. (a) Simulation results generated by the CMASE framework. (b) Predictions using the EVI index from the original study. (c) Predictions using the NDVI index from the original study. (d) Movement trajectories of agents during CMASE simulation. (e) Regional variation in agents’ emotional states.
@@
4.4. Qualitative Fieldwork
In addition to the quantitative replication, we entered the CMASE environment as one of the agents and lived alongside the others. Through CMASE, we are able to deepen the conclusions and findings of the original study via observation and interview. We conducted interview analyses with two agents who remained indoors throughout the experiment (Figure 2 d-e). Full interview transcripts and detailed analyses are provided in Appendix A. The results indicate that agents with different occupations exhibited varying mechanisms for mitigating the sense of social fragmentation in isolation. Agents in creative and analytical professions tended to rely more on orderly and low-stimulus environments, using these conditions to construct an internally stable and distinct mental world that anchored their cognitive states. For these agents, the green space had a relatively limited effect in alleviating their sense of social fragmentation.
@@ 4.4. Qualitative Fieldwork
From a language-philosophical perspective, linguistic structures and experiential structures may exhibit a certain degree of partial isomorphism or structural correspondence (Sun, 2024; Haspelmath, 2008; Pederson et al., 1998). Thus, LLMs trained on human language may partially acquire patterns of human experience (although this does not imply that they possess or actually experience these phenomena). While we remain cautious in interpreting this “embodied-like” tendency, as it could also stem from hallucinatory processes, it is equally reasonable that agents, by design, prioritize coherence with their predefined human personas. For a human persona, embodied reflections are integral to an authentic experience; thus, simulating these expressions is a necessary feature of a robust social simulation. In any case, these patterns suggest that CMASE provides emerging evidence that agents can transcend the lack of embodiment and simulate experiential phenomena, thereby offering a reliable simulation platform.
4.5. Evaluation on Agent Persona and Ethnographic Affordances
To further validate the utility of CMASE, we conducted a systematic evaluation using LLM-as-a-Judge based on the interaction logs and interview transcripts from our community isolation experiment. The resulting data are categorized into two primary areas: Agent Persona Fidelity (Table 2) and Ethnographic Affordances (Table 3). To reflect the demographic variables integrated into the agent personas within the simulated field described in section 4.1, we categorized the agents into four distinct clusters: Creative and Design (e.g., Fashion Designer), Analytical and Academic (e.g., Lawyer), Technical and Strategic (e.g., Software Engineer), and Public and Healthcare (e.g., Cardiologist).
Our assessment of agent persona quality focused on four dimensions: consistency (alignment with long-term self-awareness), responsiveness (logical feedback to environmental stimuli), and emotional alignment (accuracy of emotional state transitions). For ethnographic affordances, we measured thick description (richness of the linguistic world), traceability (clarity of the decision chain), interaction depth (sense of meaningful dialogue), and mechanistic insight (ability to reveal social logic). To ensure evaluation efficiency, we utilized Gemini-3.1 pro to evaluate the outputs across these dimensions and provide ratings within the range of [1, 5].
Table 2. Evaluation results on Agent Persona Fidelity
Persona Cluster Consistency Responsiveness Emotional Alignment Overall
Creative and Design 4.00 3.33 4.67 4.00
Analytical and Academic 4.33 3.33 4.33 4.00
Technical and Strategic 4.00 4.00 5.00 4.33
Public and Healthcare 5.00 4.50 5.00 4.83
Average 4.30 3.70 4.70 4.23
Table 3. Evaluation results on Ethnographic Affordances
Persona Cluster
Thick
Description
Traceability
Interaction
Depth
Mechanistic
Insight
Overall
Creative and Design 4.33 4.00 3.33 3.67 3.83
Analytical and Academic 5.00 4.00 4.33 4.00 4.33
Technical and Strategic 4.00 4.00 4.50 4.50 4.25
Public and Healthcare 5.00 5.00 4.50 4.50 4.75
Average 4.60 4.20 4.10 4.10 4.25
We observe that regarding agent persona fidelity, CMASE demonstrates strong performance across all metrics, with an average overall score of 4.2. Particularly impressive is the system’s capability in emotional alignment (average 4.70), indicating that the discretization method based on the VAD lexicon effectively constrains the agents’ emotional reasoning. The system also excels in consistency (4.30), ensuring that agents maintain stable self-awareness during long-term operation. The slightly lower score in responsiveness (3.70) suggests that agents’ real-time decision-making logic requires further optimization when handling high-intensity environmental feedback.
Regarding the quality of ethnographic affordances, CMASE performs prominently in thick description (average 4.60) and traceability (average 4.20). This suggests that the system effectively transforms simulated environments into virtual ethnographic fields, providing researchers with computational observation venues endowed with interpretative depth. Comparatively, while interaction depth (4.10) and mechanistic insight (4.10) are at a commendable level, there remains room for improvement. Overall, these results confirm that CMASE provides a highly interpretable and interactive foundation for computational sociology.
4.6. Time and Cost Analysis